A burglar’s car had a start with an accelaration of 2 m/s^2. A police vigilant party came in a car to the spot at a velocity of 20 m/s after 3.75 sec and continued to chase the bulgar’s car with uniform velocity. Find the time in which the police van  will overtake the car.

Grade:11

2 Answers

Carmel Emmanuel
8 years ago
Distance covered by burglar in 3.75
 
S=1/2at2 =14.06
Velocity of burglar at that moment=at=7.50
At this time police arrived 
Let t be time to overtake
urelative=20-7.5=12.5
Now apply equation of motion 
 
Sf-Si=urelativet+1/2arelativet2
sf=0 Si=14.06 
Find out the roots of t
Priya
5 years ago
u=0
a=2
S = ut+1/2at^2
20t= 0(t+5)+ 1/2×2(t+5)^2
20t=(t+5)^2
20t=t^2+25+10t
t^2-10t+25=0
(t-5)^2=0
t-5=0
t=5sec